OpenStudy (richyw):

Need help with a trigonometric identity. I have \[E_y=E_{0y}\cos (kz-\omega t+\epsilon)\]and my book says that this is equivalent to \[\frac{E_y}{E_{0y}} =\cos (kz-\omega t)\cos (\epsilon) -\sin (kz-\omega t)\sin (\epsilon)\]. Would someone be able to let me know how my book derived this identity?

OpenStudy (nincompoop):

it's not rendering properly

OpenStudy (richyw):

fixed

OpenStudy (richyw):

also nevermind. I just figure out it's just \[cos (a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\]right?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

:D